OpenStudy (anonymous):

Sixteen squared divided by four to the fourth power, write as a given expression and then evaluate.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

1.....the answer is 1.

OpenStudy (anonymous):

16^2/4^4 is what you wrote

OpenStudy (anonymous):

and yes the answer is 1 when you solve for it because 16*16=256 and 4*4*4*4=256 256/256=1